one hundred eight americans were surveyed determine. the number of hours they spend watching tv each month. it was revealed that they watched an average of 151hours each month with a standard deviation of 32 hours. Assume that the underlying population destitution is normal.

b. find critical value of 99% population mean time spent waiting

c. which distribution should you use for this problem?

d.calculate error bound

e. sketch graph

f. what are the lower and upper bounds
Answer #1

Solution :

Given that,

Point estimate = sample mean = = 151  


Population standard deviation = s = 32

Sample size = n = 108

b) At 99% confidence level

= 1-0.99% =1-0.99 =0.01

/2 =0.01/ 2= 0.005

t/2,df = t0.005,107 = 2.62

critical value = t/2,df = 2.62

c)singal sampla t value confidence interval useing

d) Margin of error = E = t/2,df * (s /n)

= 2.62 * (32 / 108)

Margin of error = E = 8.08

The 99% confidence interval estimate of the population mean is,

- E < <  + E

151 - 8.08 < < 151 + 8.08

142.92 < < 159.08

(142.92 ,159.08)

e)

f)

lower bound = 142.92

upper bound = 159.08